# The inducibility of Turán graphs

### v1

@@ 4.1 Preparations
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph on $\ell\in[r+1,~2r-1]$ vertices (i.e. $\ell=a_{1}+\cdots+a_{r}$). Let $m=m_{r,\ell}$ be the integer given by Lemma 3.2. Assume that $a_{1}\geq\cdots\geq a_{r}$. By the Pigeonhole Principle, we have $a_{r}=1$. By the definition of almost balanced (see (1)), we have $\binom{a_{1}-1}{2}<1$, which means that $a_{1}\leq 2$. Thus we have
It follows from Theorem 3.1 that $\mathrm{OPT}(F)=\{\bm{m}\}$. Thus, by Theorem 4.1, it suffices to verify (S1) and (S2) with some constant $\varepsilon>0$ and all sufficiently large $n$.
Fix a sufficiently large $n$. For simplicity, we will use $o(\cdot)$ to denote lower-order terms and assume that $n=km$ for some integer $k$ (i.e., $n$ is divisible by $m$). It will be clear from the proof that this divisibility assumption has no substantive effect; it is included only to allow us to omit floor and ceiling signs.
Let $G\coloneqq G_{n,\bm{m}}$ be the $n$-vertex realization with parts $V_{1},\ldots,V_{m}$, where $|V_{i}|=k$ for all $i\in[m]$. Let $\varepsilon>0$ be the constant satisfying
$1/\varepsilon=m^{\ell+1}{2^{\ell-r}(\ell-r)!(2r-\ell)!}.$
@@
By symmetry, we may assume that $(x,y)\in V_{1}\times V_{2}$. Since $x$ and $y$ lie in different parts, the pair $\{x,y\}$ is an edge of $G$.
Every induced copy $T$ of $F$ in $G$ that contains the edge $xy$ can be classified into three cases, depending on how its remaining vertices are distributed among the parts $V_{1},\ldots,V_{m}$}.
Case 1: $|V_{1}\cap V(T)|=|V_{2}\cap V(T)|=1$.
If $\ell=2r-1$, the number of the copies of these $T$ is 0, otherwise this number is
$\frac{(m-2)_{r-2}\cdot k^{2r-\ell-2}\cdot\binom{k}{2}^{\ell-r}}{(\ell-r)!(2r-\ell-2)!}=\frac{(m-2)_{r-2}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)}{2^{\ell-r}(\ell-r)!(2r-\ell-2)!}.$ (25)
To see this, note that $x$ and $y$ are the unique vertices from $V_{1}$ and $V_{2}$ in $T$. The factor $(m-2)_{r-2}$ indicates the number of ways to choose the r2r-2 other parts of $T$ from the remaining m2m-2 parts (of size $k$). Among the r2r-2 other parts of $T$, there are $2r-\ell-2$ parts of size $1$ and $\ell-r$ parts of size $2$, so the denominator $(\ell-r)!(2r-\ell-2)!$ accounts for overcounting due to permutations of parts of equal size. If $\ell=2r-1$, no part apart from $V_{1},V_{2}$ can have size $1$, so the count is 0.
Using similar estimations, we can obtain the remaining two cases.
Case 2: $|V_{1}\cap V(T)|=1,|V_{2}\cap V(T)|=2$ or $|V_{1}\cap V(T)|=2,|V_{2}\cap V(T)|=1$.
The number of the induced copies of these $T$ is
$2\cdot\frac{(m-2)_{r-2}\cdot(k-1)k^{2r-\ell-1}\cdot\binom{k}{2}^{\ell-r-1}}{(\ell-r-1)!(2r-\ell-1)!}=\frac{(m-2)_{r-2}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)}{2^{\ell-r-2}(\ell-r-1)!(2r-\ell-1)!}.$
Case 3: $|V_{1}\cap V(T)|=|V_{2}\cap V(T)|=2$. If $\ell=r+1$, the number of the induced copies of these $T$ is 0, otherwise this number is
$\frac{(m-2)_{r-2}\cdot(k-1)^{2}k^{2r-\ell}\cdot\binom{k}{2}^{\ell-r-2}}{(\ell-r-2)!(2r-\ell)!}=\frac{(m-2)_{r-2}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)}{2^{\ell-r-2}(\ell-r-2)!(2r-\ell)!}.$
Summing over all cases (regardless of the relative sizes of $\ell$ and $r$), the number of $T$ containing $xy$ is
$\frac{(m-2)_{r-2}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\cdot\bigl((2r-\ell)(2r-\ell-1)+4(\ell-r)(2r-\ell)+4(\ell-r)(\ell-r-1)\bigr)$
$={}$ $\frac{(m-2)_{r-2}(\ell^{2}+2r-3\ell)}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr).$
Now we turn to the induced copy $T$ of $F$ in $G\oplus xy$ containing non-edge $xy$, so $x,y$ must constitute a part of $T$. Then the number of the induced copies of these $T$ is
$\frac{(m-2)_{r-1}k^{2r-\ell}\binom{k}{2}^{\ell-r-1}}{(\ell-r-1)!(2r-\ell)!}=\frac{(m-2)_{r-1}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)}{2^{\ell-r-1}(\ell-r-1)!(2r-\ell)!}.$
Hence we obtain
$I(F,G)-I(F,G\oplus xy)=$ $\frac{(m-2)_{r-2}(\ell^{2}+2r-3\ell-2(m-r)(\ell-r))}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)$
$=$ $\frac{(m-2)_{r-2}(\ell^{2}+2r-3\ell-2(m-r)(\ell-r))}{2^{\ell-r}m^{\ell-2}(\ell-r)!(2r-\ell)!}\bigl(n^{\ell-2}+o(n^{\ell-2})\bigr).$
By Lemma 4.2 (ii), we have $m<\frac{\ell(\ell-1)}{2(\ell-r)}$. Substituting this upper bound gives
$\ell^{2}+2r-3\ell-2(m-r)(\ell-r)$ $>\ell^{2}+2r-3\ell-2\left(\frac{\ell(\ell-1)}{2(\ell-r)}-r\right)(\ell-r)$
$=\ell^{2}+2r-3\ell-\ell(\ell-1)+2r(\ell-r)$
$=2(r-1)(\ell-r)\geq 2.$
Hence $I(F,G)-I(F,G\oplus xy)>\varepsilon n^{\ell-2}$ for all sufficiently large $n$. This completes the proof of Claim 4.4. ∎
@@
Assume without loss of generality that $x,y\in V_{1}$ are two distinct vertices. By the definition of $G$, the edge $xy$ is absent from $G$.
The number of induced copies of $F$ in $G$ that contain the non‑edge $xy$ is
$\frac{(m-1)_{r-1}k^{2r-\ell}\binom{k}{2}^{\ell-r-1}}{(\ell-r-1)!(2r-\ell)!}=\frac{(m-1)_{r-1}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)}{2^{\ell-r-1}(\ell-r-1)!(2r-\ell)!}.$
If $\ell=2r-1$, then $F$ has no part of size larger than $1$ apart from the two vertices $x,y$; consequently, there is no induced copy of $F$ in $G\oplus xy$ that contains the edge $xy$. Otherwise, the number of such copies is
$\frac{(m-1)_{r-2}k^{2r-\ell-2}\binom{k}{2}^{\ell-r}}{(\ell-r)!(2r-\ell-2)!}=\frac{(m-1)_{r-2}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)}{2^{\ell-r}(\ell-r)!(2r-\ell-2)!}.$
Therefore, regardless of the relation between $\ell$ and $r$, we have
$I(F,G)-I(F,G\oplus xy)$ $=\frac{(m-1)_{r-2}\bigl(2(\ell-r)(m-r+1)-(2r-\ell)(2r-\ell-1)\bigr)}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)$
$=\frac{(m-1)_{r-2}\bigl(2(\ell-r)m+\ell-\ell^{2}+2r\ell-2r^{2}\bigr)}{2^{\ell-r}m^{\ell-2}(\ell-r)!(2r-\ell)!}\bigl(n^{\ell-2}+o(n^{\ell-2})\bigr).$
By Lemma 4.2 (i), $m>\frac{\ell(r-1)}{2(\ell-r)}$. Substituting this lower bound gives
$2(\ell-r)m+\ell-\ell^{2}+2r\ell-2r^{2}>\ell(r-1)+\ell-\ell^{2}+2r\ell-2r^{2}=(\ell-r)(2r-\ell).$
Since $\ell-r$ and $2r$-$\ell$ are positive integers, their product is at least 11. Hence for all sufficiently large $n$,
$I(F,G)-I(F,G\oplus xy)>\varepsilon n^{\ell-2}.$
This completes the proof of Claim 4.5. ∎
Combining Claim 4.4 and Claim 4.5, (S1) holds. Now we turn to verify (S2).
@@
For every set $A\subseteq[m]$ with $|A|\neq m-1$,
$I(F,G_{A^{*}})-I(F,G_{A})\geq\varepsilon n^{\ell-1},$
where $A^{*}$ is any subset of $[m]$ with $|A^{*}|=m-1$.
@@
For any subset $B\subseteq[m]$ with $|B|=q$, we proceed to compute $I(F,G_{B})-I(F,G)$. Note that the induced copies (say $T$) of $F$ in $G_{B}$ containing the new vertex $v_{B}$ fall into two categories:
Case 1: $v_{B}$ is adjacent to every other vertex of $T$. If q<r1q<r-1, the number of the induced copies of these $T$ is 0; otherwise this number is
$\frac{(q)_{r-1}\cdot k^{2r-\ell-1}\cdot\binom{k}{2}^{\ell-r}}{(\ell-r)!(2r-\ell-1)!}=\frac{(q)_{r-1}\bigl(k^{\ell-1}+o(k^{\ell-1})\bigr)}{2^{\ell-r}(\ell-r)!(2r-\ell-1)!}.$
Case 2: $v_{B}$ is not adjacent to every other vertex of $T$. If q<r1q<r-1, the number of the induced copies of these $T$ is 0; otherwise this number is
$\frac{(q)_{r-1}\cdot(m-q)\cdot k^{2r-\ell+1}\cdot\binom{k}{2}^{\ell-r-1}}{(\ell-r-1)!(2r-\ell)!}=\frac{(m-q)(q)_{r-1}\bigl(k^{\ell-1}+o(k^{\ell-1})\bigr)}{2^{\ell-r-1}(\ell-r-1)!(2r-\ell)!}.$
Adding these two cases together, regardless of the value of $q$, we obtain
$I(F,G_{B})-I(F,G)=\frac{(q)_{r-1}\cdot\bigl(2r-\ell+2(m-q)(\ell-r)\bigr)\cdot\bigl(k^{\ell-1}+o(k^{\ell-1})\bigr)}{2^{\ell-r}(\ell-r)!(2r-\ell)!}.$ (26)
Define for integers $q\in[r-1,m]$,
$h(q)\coloneqq(q)_{r-1}\cdot\bigl(2r-\ell+2(m-q)(\ell-r)\bigr).$ (27)
For convenience, we extend $h$ by setting $h(q)=0$ for all integers q<r1q<r-1. We will show that for all integers $q\in[m]\setminus\{m-1\}$,
$h(q)<h(m-1).$ (28)
Since $h(m-1)=\ell(m-1)_{r-1}$, it suffices to prove that for every integer $q\in[r-1,m]\setminus\{m-1\}$,
$(q)_{r-1}\cdot\bigl(2r-\ell+2(m-q)(\ell-r)\bigr)<\ell\cdot(m-1)_{r-1}.$ (29)
For $q=m$, inequality (29) becomes $(m)_{r-1}(2r-\ell)<\ell(m-1)_{r-1}$, which simplifies to $2(\ell-r)m>\ell(r-1)$. This is precisely the lower bound on $m$ given by Lemma 4.2 (i), and hence this case holds.
We now prove (29) by reverse induction on $q$, starting from q=m2q=m-2 down to q=r1q=r-1.
For the base case q=m2q=m-2, we need to show $(m-2)_{r-1}(3\ell-2r)<\ell(m-1)_{r-1}$, which is equivalent to $2(\ell-r)m<3r\ell-2r^{2}$-$\ell$. By the upper bound in Lemma 4.2 (ii), we have $m<\frac{\ell(\ell-1)}{2(\ell-r)}$; consequently,
$2(\ell-r)m-(3r\ell-2r^{2}-\ell)<\ell(\ell-1)-3r\ell+2r^{2}+\ell=(\ell-r)(\ell-2r)<0,$
where the last inequality uses $r<\ell<2r$. Thus the base case holds.
Now we turn to inductive step. Assume that (29) holds for some integer $q$ with $r\leq q\leq m-2$. We now prove it for q1q-1. Using the inductive hypothesis, we have
$\ \ell(m-1)_{r-1}-(q-1)_{r-1}\bigl(2r-\ell+2(m-q+1)(\ell-r)\bigr)$
$>$ $\ (q)_{r-1}\bigl(2r-\ell+2(m-q)(\ell-r)\bigr)-(q-1)_{r-1}\bigl(2r-\ell+2(m-q+1)(\ell-r)\bigr)$
$=$ $\ (q-1)_{r-2}\Bigl(2(\ell-r)\bigl((r-1)(m+1)-rq\bigr)+(r-1)(2r-\ell)\Bigr)$
$\geq$ $\ (q-1)_{r-2}\Bigl(2(\ell-r)(3r-m-1)+(r-1)(2r-\ell)\Bigr),$
where the last inequality follows from the assumption $q\leq m-2$. Indeed,
$(r-1)(m+1)-rq\geq(r-1)(m+1)-r(m-2)=3r-m-1.$
Using the upper bound $m<\frac{\ell(\ell-1)}{2(\ell-r)}$ from Lemma 4.2 (ii), we further have
$2(\ell-r)(3r-m-1)+(r-1)(2r-\ell)$ $>2(\ell-r)\left(3r-1-\frac{\ell(\ell-1)}{2(\ell-r)}\right)+(r-1)(2r-\ell)$
$=2(\ell-r)(3r-1)-\ell(\ell-1)+(r-1)(2r-\ell)$
$=5r\ell-\ell^{2}-4r^{2}=(\ell-r)(4r-\ell)>0.$
This proves that $\ell(m-1)_{r-1}-(q-1)_{r-1}\bigl(2r-\ell+2(m-q+1)(\ell-r)\bigr)>0,$ which is precisely (29) for q1q-1. Thus (28) holds.
Combining (26), (27), and (28), we obtain that for all sufficiently large $n$ and for any $A,A^{*}\subseteq[m]$ with $|A^{*}|=m-1$ and $|A|=p\neq m-1$,
$I(F,G_{A^{\ast}})-I(F,G_{A})$ $=\bigl(I(F,G_{A^{\ast}})-I(F,G)\bigr)-\bigl(I(F,G_{A})-I(F,G)\bigr)$
$=\bigl(h(m-1)-h(p)\bigr)\cdot\frac{k^{\ell-1}+o(k^{\ell-1})}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\geq\frac{k^{\ell-1}+o(k^{\ell-1})}{2^{\ell-r}(\ell-r)!(2r-\ell)!}$
$=\frac{n^{\ell-1}+o(n^{\ell-1})}{2^{\ell-r}m^{\ell-1}(\ell-r)!(2r-\ell)!}>\varepsilon n^{\ell-1}.$
Hence, Claim 4.6 holds. ∎
By Theorem 4.1, all almost balanced complete $r$-partite graphs on $\ell\in[r+1,~2r-1]$ vertices is perfectly stable. This completes the proof of Theorem 1.5 for $\ell\leq 2r-1$. ∎
@@
In this subsection, we prove Theorem 1.5 for the case $\ell\geq 2r$. The proof parallels that of the previous section, with several steps simplified by the fact that $\ell\geq 2r$ implies $a_{r}\geq 2$.

### v2

@@ 4.1 Preparations
Lemma 4.4.
Let $\ell,r$ be integers such that $r\geq 2$ and $r+1\leq\ell\leq 2r-1$. Let $m=m_{r,\ell}$ be the integer given by Lemma 3.2. Then the discrete function $h\colon[0,m]\to\mathbb{Z}$ defined by
$h(q)\coloneqq(q)_{r-1}\cdot\bigl(2r-\ell+2(m-q)(\ell-r)\bigr),\quad\text{for every}\quad q\in[0,m],$
satisfies h(q)<h(m1)h(q)<h(m-1) for all integers $q\in[0,m]\setminus\{m-1\}$.
Proof.
Note that $h(m-1)=\ell(m-1)_{r-1}>0$. So it suffices to prove that for every $q\in[r-1,m]\setminus\{m-1\}$,
$h(q)=(q)_{r-1}\cdot\bigl(2r-\ell+2(m-q)(\ell-r)\bigr)<\ell(m-1)_{r-1}.$ (26)
For $q=m$, inequality (26) becomes $(m)_{r-1}(2r-\ell)<\ell(m-1)_{r-1}$, which simplifies to $2(\ell-r)m>\ell(r-1)$. This follows from the lower bound on $m$ given by Lemma 4.2 (i), and hence this case holds. It remains to show h(q)<h(m1)h(q)<h(m-1) for every $q\in[r,m-2]$.
Claim 4.5.
Suppose that $q\in[r,m-1]$. Then h(q)h(q1)>0h(q)-h(q-1)>0.
Proof.
First, we consider the case q=m1q=m-1. In this case we need to show $(m-2)_{r-1}(3\ell-2r)<\ell(m-1)_{r-1}$, which is equivalent to $2(\ell-r)m<3r\ell-2r^{2}$-$\ell$. Since $m<\frac{\ell(\ell-1)}{2(\ell-r)}$ (by Lemma 4.2 (ii)), we have
$2(\ell-r)m-(3r\ell-2r^{2}-\ell)<2(\ell-r)\cdot\frac{\ell(\ell-1)}{2(\ell-r)}-(3r\ell-2r^{2}-\ell)=(\ell-r)(\ell-2r)<0.$
This proves that h(m2)<h(m1)h(m-2)<h(m-1).
Next, we consider the case $q\in[r,m-2]$. It follows from $m<\frac{\ell(\ell-1)}{2(\ell-r)}$ (by Lemma 4.2 (ii)) that
$2(\ell-r)(3r-m-1)+(r-1)(2r-\ell)$ $>2(\ell-r)\left(3r-1-\tfrac{\ell(\ell-1)}{2(\ell-r)}\right)+(r-1)(2r-\ell)$
$=2(\ell-r)(3r-1)-\ell(\ell-1)+(r-1)(2r-\ell)$
$=(\ell-r)(4r-\ell)>0.$
Since $q\leq m-2$, we have
$(r-1)(m+1)-rq\geq(r-1)(m+1)-r(m-2)=3r-1-m.$
Consequently,
$2(\ell-r)\bigl((r-1)(m+1)-rq\bigr)+(r-1)(2r-\ell)\geq 2(\ell-r)(3r-m-1)+(r-1)(2r-\ell)>0.$ (27)
Now, fix $q\in[r,m-2]$, and let $\delta(q)\coloneqq h(q)-h(q-1)$. Then
$\delta(q)$ $=(q)_{r-1}\bigl(2r-\ell+2(m-q)(\ell-r)\bigr)-(q-1)_{r-1}\bigl(2r-\ell+2(m-q+1)(\ell-r)\bigr)$
$=(q-1)_{r-2}\left(q\bigl(2r-\ell+2(m-q)(\ell-r)\bigr)-(q-r+1)\bigl(2r-\ell+2(m-q+1)(\ell-r)\bigr)\right)$
$=(q-1)_{r-2}\left(2(\ell-r)\bigl((r-1)(m+1)-rq\bigr)+(r-1)(2r-\ell)\right)>0,$
where the last inequality follows from (27). This proves Claim 4.5. ∎
It follows from Claim 4.5 that h(q)<h(m1)h(q)<h(m-1) for every $q\in[r,m-2]$. This proves Lemma 4.4. ∎
@@
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph on $\ell\in[r+1,~2r-1]$ vertices (i.e. $\ell=a_{1}+\cdots+a_{r}$). Let $m=m_{r,\ell}$ be the integer given by Lemma 3.2. Assume that $a_{1}\geq\cdots\geq a_{r}$. By Fact 2.7 (i), we have $a_{1}=\cdots=a_{\ell-r}=2$ and $a_{\ell-r+1}=\cdots=a_{r}=1$.
Let $\varepsilon>0$ be a sufficiently small constant such that, in particular,
$\varepsilon\leq\left(2m^{\ell+1}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\right)^{-1}.$
It follows from Theorem 3.1 that $\mathrm{OPT}(F)=\{\bm{m}\}$. Thus, by Theorem 4.1, it suffices to verify (S1) and (S2) for all sufficiently large $n$ (with $\varepsilon$ there corresponding to $\varepsilon$ here).
Fix a sufficiently large integer $n$ and let $k\coloneqq\lfloor n/m\rfloor$. For convenience, we will use $o(\cdot)$ to denote lower-order terms. Let $G\coloneqq G_{n,\bm{m}}$ be the $n$-vertex realization with parts $V_{1},\ldots,V_{m}$, where $|V_{i}|=k$ for all $i\in[m]$.
@@
By symmetry, we may assume that $(x,y)\in V_{1}\times V_{2}$. Since $x$ and $y$ lie in different parts, the pair $\{x,y\}$ is an edge of $G$. Let $\tilde{G}$ denote the induced subgraph of $G$ on $V_{3}\cup\cdots\cup V_{r}$, noting that $\tilde{G}$ is complete (m2)(m-2)-partite.
Let $\mathcal{C}$ denote the collection of all induced copies of $F$ in $G$ that contain $\{x,y\}$. Let
$\mathcal{C}_{1,1}$ $\coloneqq\left\{T\in\mathcal{C}\colon\left(|V_{1}\cap V(T)|,|V_{2}\cap V(T)|\right)=(1,1)\right\},$
$\mathcal{C}_{1,2}$ $\coloneqq\left\{T\in\mathcal{C}\colon\left(|V_{1}\cap V(T)|,|V_{2}\cap V(T)|\right)=(1,2)\right\},$
$\mathcal{C}_{2,1}$ $\coloneqq\left\{T\in\mathcal{C}\colon\left(|V_{1}\cap V(T)|,|V_{2}\cap V(T)|\right)=(2,1)\right\},$
$\mathcal{C}_{2,2}$ $\coloneqq\left\{T\in\mathcal{C}\colon\left(|V_{1}\cap V(T)|,|V_{2}\cap V(T)|\right)=(2,2)\right\},$
noting that $\mathcal{C}_{1,1}\cup\mathcal{C}_{1,2}\cup\mathcal{C}_{2,1}\cup\mathcal{C}_{2,2}=\mathcal{C}$.
[…]
$|\mathcal{C}_{1,1}|$ $=\begin{cases}0,&\quad\text{if}\quad\ell=2r-1\\[1.99997pt] \frac{(1+o(1))(m-2)_{r-2}}{2^{\ell-r}(\ell-r)!(2r-\ell-2)!}k^{\ell-2}&\quad\text{if}\quad\ell\leq 2r-2\end{cases}\quad=\frac{(1+o(1))(m-2)_{r-2}}{2^{\ell-r}(\ell-r)!(2r-\ell)!}(2r-\ell)(2r-1-\ell)k^{\ell-2}.$
Similar, we have
$|\mathcal{C}_{1,2}|+|\mathcal{C}_{1,2}|$ $=2\cdot I\big(K_{a_{2},\ldots,a_{r-1}},\tilde{G}\big)=\frac{(1+o(1))(m-2)_{r-2}}{2^{\ell-r-2}(\ell-r-1)!(2r-\ell-1)!}k^{\ell-2},$
and
$|\mathcal{C}_{2,2}|=\begin{cases}0&\text{if}\quad\ell=r+1\\[1.99997pt] \frac{(1+o(1))(m-2)_{r-2}}{2^{\ell-r-2}(\ell-r-2)!(2r-\ell)!}k^{\ell-2}&\text{if}\quad\ell\geq r+2\end{cases}\quad=\frac{(1+o(1))(m-2)_{r-2}}{2^{\ell-r-2}(\ell-r)!(2r-\ell)!}(\ell-r)(\ell-r-1)k^{\ell-2}.$
Therefore,
$|\mathcal{C}|$ $=|\mathcal{C}_{1,1}|+|\mathcal{C}_{1,2}|+|\mathcal{C}_{1,2}|+|\mathcal{C}_{2,2}|$
$=\frac{(1+o(1))(m-2)_{r-2}}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\left((2r-\ell)(2r-1-\ell)+4(\ell-r)(2r-\ell)+4(\ell-r)(\ell-r-1)\right)k^{\ell-2}$
$=(1+o(1))\frac{(m-2)_{r-2}(\ell^{2}-3\ell+2r)}{2^{\ell-r}(\ell-r)!(2r-\ell)!}k^{\ell-2}.$ (28)
[…]
$|\mathcal{C}^{\prime}|=\frac{(1+o(1))(m-2)_{r-1}}{2^{\ell-r-1}(\ell-r-1)!(2r-\ell)!}k^{\ell-2}.$
Combining this with (4.2), we obtain
$I(F,G)-I(F,G\oplus xy)=|\mathcal{C}|-|\mathcal{C}^{\prime}|$ $=\frac{(1+o(1))(m-2)_{r-2}}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\big(\ell^{2}-3\ell+2r-2(m-r)(\ell-r)\big)k^{\ell-2}.$
By Lemma 4.2 (ii), we have $m<\frac{\ell(\ell-1)}{2(\ell-r)}$, which implies that
$\ell^{2}-3\ell+2r-2(m-r)(\ell-r)$ $>\ell^{2}-3\ell+2r-2\left(\tfrac{\ell(\ell-1)}{2(\ell-r)}-r\right)(\ell-r)$
$=\ell^{2}-3\ell+2r-\ell(\ell-1)+2r(\ell-r)=2(r-1)(\ell-r).$
Therefore,
$I(F,G)-I(F,G\oplus xy)\geq(1+o(1))\frac{2(r-1)(\ell-r)(m-2)_{r-2}}{2^{\ell-r}(\ell-r)!(2r-\ell)!}k^{\ell-2}>\varepsilon n^{\ell-2},$
which proves Claim 4.6. ∎
@@
By symmetry, we may assume that $\{x,y\}\subseteq V_{1}$. Note that the pair $\{x,y\}$ is not an edge of $G$. Let $G^{\prime}$ denote the induced subgraph of $G$ on $V_{2}\cup\cdots\cup V_{m}$.
It is clear that the number of induced copies of $F$ in $G$ that contain $\{x,y\}$ coincides with the value of $I(K_{a_{2},\ldots,a_{r}},G^{\prime})$, which, by Fact 2.2, is
$(1+o(1))\frac{(m-1)_{r-1}}{2^{\ell-r-1}(\ell-r-1)!(2r-\ell)!}k^{\ell-2}.$
If $\ell=2r-1$, then $F$ has only one part of size one, and hence, the number of induced copies of $F$ in $G\oplus xy$ that contain $\{x,y\}$ is 0. If $\ell\leq 2r-2$, then the number of induced copies of $F$ in $G\oplus xy$ that contain $\{x,y\}$ coincides with the value of $I(K_{a_{1},\ldots,a_{r-2}},G^{\prime})$, which, by Fact 2.2, is
$(1+o(1))\frac{(m-1)_{r-2}}{2^{\ell-r}(\ell-r)!(2r-\ell-2)!}k^{\ell-2}.$
Therefore,
$I(F,G)-I(F,G\oplus xy)$ $=\frac{(1+o(1))(m-1)_{r-2}}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\big(2(\ell-r)(m-r+1)-(2r-\ell)(2r-1-\ell)\big)k^{\ell-2}.$
By Lemma 4.2 (i), we have $m>\frac{\ell(r-1)}{2(\ell-r)}$. Therefore,
$2(\ell-r)(m-r+1)-(2r-\ell)(2r-1-\ell)\\ >2(\ell-r)\left(\tfrac{\ell(r-1)}{2(\ell-r)}-r+1\right)-(2r-\ell)(2r-1-\ell)=(\ell-r)(2r-\ell).$
It follows that
$I(F,G)-I(F,G\oplus xy)$ $\geq\frac{(1+o(1))(m-1)_{r-2}}{2^{\ell-r}(\ell-r)!(2r-\ell)!}(\ell-r)(2r-\ell)k^{\ell-2}$
$=\frac{(1+o(1))(m-1)_{r-2}}{2^{\ell-r}(\ell-r-1)!(2r-\ell-1)!}k^{\ell-2}>\varepsilon n^{\ell-2},$
which proves Claim 4.7. ∎
It follows from Claims 4.6 and 4.7 that (S1) holds. So it remains to verify (S2).
@@
Suppose that $B\subseteq[m]$ is a subset with $|B|=q\in[0,m]$. Then
$|\tilde{\mathcal{C}}_{q}|$ $=\frac{(1+o(1))h(q)}{2^{\ell-r}(\ell-r)!(2r-\ell)!}k^{\ell-1},$
where $h\colon[m]\to\mathbb{Z}$ is the discrete function defined in Lemma 4.4.
@@
Fix a subset $B\subseteq[m]$ with $|B|=q$. Recall that $G_{B}$ is the graph obtained from $G$ by adding a new vertex $v_{\ast}$ that is adjacent to precisely all vertices in $\bigcup_{i\in B}V_{i}$. By symmetry, we may assume that $B=\{1,\ldots,q\}$.
Let $\tilde{\mathcal{C}}_{q}$ denote the collection of all induced copies of $F$ in $G_{A}$ that contain $v_{\ast}$. Let $\tilde{\mathcal{C}}_{q}^{1}$ denote the collection of all induced copies of $F$ in $G_{A}$ in which $v_{\ast}$ forms a part of size one, and let $\tilde{\mathcal{C}}_{q}^{2}$ denote the collection of all induced copies of $F$ in $G_{A}$ in which $v_{\ast}$ lies in a part of size two.
Note that the size of $\tilde{\mathcal{C}}_{q}^{1}$ coincides with the value of $I(K_{a_{1},\ldots,a_{r-1}},K[V_{1},\ldots,V_{q}])$, which, by Fact 2.2, is
$\frac{(1+o(1))(q)_{r-1}}{2^{\ell-r}(\ell-r)!(2r-\ell-1)!}k^{\ell-1}.$
Note that the size of $\tilde{\mathcal{C}}_{q}^{2}$ coincides with the value of $I(K_{a_{2},\ldots,a_{r}},K[V_{1},\ldots,V_{q})\cdot\sum_{i=q+1}^{m}|V_{i}|$, which, by Fact 2.2, is
$\frac{(1+o(1))(q)_{r-1}}{2^{\ell-r-1}(\ell-r-1)!(2r-\ell)!}k^{\ell-2}\cdot(m-q)k=\frac{(1+o(1))(m-q)(q)_{r-1}}{2^{\ell-r-1}(\ell-r-1)!(2r-\ell)!}k^{\ell-1}.$
Therefore,
$|\tilde{\mathcal{C}}_{q}|=|\tilde{\mathcal{C}}_{q}^{1}|+|\tilde{\mathcal{C}}_{q}^{2}|$ $=\frac{(1+o(1))(q)_{r-1}}{2^{\ell-r}(\ell-r)!(2r-\ell-1)!}k^{\ell-1}+\frac{(1+o(1))(m-q)(q)_{r-1}}{2^{\ell-r-1}(\ell-r-1)!(2r-\ell)!}k^{\ell-1}$
$=\frac{(1+o(1))(q)_{r-1}}{2^{\ell-r}(\ell-r)!(2r-\ell)!}\big((2r-\ell)+2(m-q)(\ell-r-1)\big)k^{\ell-1}=\frac{(1+o(1))h(q)}{2^{\ell-r}(\ell-r)!(2r-\ell)!}k^{\ell-1}.$
This proves Claim 4.8. ∎
Fix an arbitrary set $A\subseteq[m]$ with $|A|\neq m-1$. Let
$\Delta\coloneqq\min\big\{I(F,G_{A^{\ast}})\colon A^{\ast}\subseteq[m]\text{ and }|A^{\ast}|=m-1\big\}-I(F,G_{A}).$
Then it follows from Claim 4.8 and Lemma 4.4 that
$\Delta$ $=\frac{(1+o(1))h(m-1)}{2^{\ell-r}(\ell-r)!(2r-\ell)!}k^{\ell-1}-\frac{(1+o(1))h(|A|)}{2^{\ell-r}(\ell-r)!(2r-\ell)!}k^{\ell-1}$
$=\frac{h(m-1)-h(|A|)}{2^{\ell-r}(\ell-r)!(2r-\ell)!}k^{\ell-1}+o(k^{\ell-1})\geq\frac{1}{2^{\ell-r}(\ell-r)!(2r-\ell)!}k^{\ell-1}+o(k^{\ell-1})\geq\varepsilon n^{\ell-1}.$
This shows that (S2) holds. So it follows from Theorem 4.1 that the inducibility problem for $F$ is perfectly stable, which completes the proof of Theorem 1.5 for $\ell\leq 2r-1$. ∎
@@
In this subsection, we prove Theorem 1.5 for the case $\ell\geq 2r$. The proof parallels that of the previous section, with several steps simplified by the fact that $a_{r}\geq 2$ (which follows from $\ell\geq 2r$).
Lemma 4.9.
Let $\ell,r$ be integers such that $r\geq 2$ and $\ell\geq 2r$. Let $m=m_{r,\ell}$ be the integer given by Lemma 3.2. Then the discrete function $H\colon[0,m]\to\mathbb{Z}$ defined by
$H(q)\coloneqq(m-q)(q)_{r-1},\quad\text{for every}\quad q\in[0,m],$
satisfies H(q)<H(m1)H(q)<H(m-1) for all integers $q\in[0,m]\setminus\{m-1\}$.
Proof.
The inequality H(q)<H(m1)H(q)<H(m-1) is clear for $q=m$ and $q\leq r-2$ since $H(q)=0$ in these cases. For the remaining case, it suffices to show that $H(q)$ is increasing in $q$ for $q\in[r-1,m-2]$.
It follows from Lemma 4.2 (ii) that
Fix $q\in[r,m-1]$. Then we have
$H(q)-H(q-1)$ $=(m-q)(q)_{r-1}-(m-q+1)(q-1)_{r-1}$
$=(q-1)_{r-2}\big((m-q)q-(m-q+1)(q-r+1)\big)$
$=(q-1)_{r-2}\big((r-1)(m+1)-rq\big)$
$\geq(q-1)_{r-2}\big(3r-1-m\big)>(q-1)_{r-2}\left(\tfrac{5r}{3}-1\right)>0.$
Here, in the first inequality we used the assumption that $q\leq m-2$, in the second inequality we used (29). This completes the proof of Lemma 4.9. ∎